OpenStudy (anonymous):

Solve the equation. 6=2(x+8)-5x A.2/3 B.^3 1/3 C.-2/3 D.^3 1/3

OpenStudy (king):

x=10/3

OpenStudy (anonymous):

2x+16-5x=6 3x=10 x=10/3